OpenStudy (anonymous):

Look at the pattern shown below. What number is missing from the pattern? 20, 22, 26, 34, __,82

OpenStudy (anonymous):

Welcome to OpenStudy and do you see any pattern here? They're all even #'s for one ;-)

OpenStudy (anonymous):

20

OpenStudy (anonymous):

34 + 16 = ??

OpenStudy (anonymous):

@fadxi No.

OpenStudy (anonymous):

Hmm at first I thought it was \(\Sigma_1^\infty\) [ {previous term} + 2(3+n) ]

OpenStudy (anonymous):

But 82 is a big jump, so now I'm thinking it's an exponential

OpenStudy (anonymous):

@ladiic03 Can you please list out the prime factorization of each of those numbers? Like so: 22=(2)(11) 26=(2)(13) 34=? 82=?

OpenStudy (anonymous):

That's the trick if you're looking for an exponential

OpenStudy (anonymous):

HAHA GOT IT! :-D

OpenStudy (anonymous):

You still there @ladiic03?

OpenStudy (anonymous):

\[\large 2(\frac{a_{n-1}}{2}+2^n),a_0=11,a_1\ to\ a_4\]

OpenStudy (anonymous):

22 / 2 = 11 26 / 2 = 13 34 / 2 = 17 50 / 2 = 25 82 / 2 = 41 13 - 11 = 2 17 - 13 = 4 25 - 17 = 8 41 - 25 = 16 2^1=2 2^2=4 2^3=8 2^4=16 ;-)